OpenStudy (anonymous):

Subtract -1/35- (-9/40)

OpenStudy (anonymous):

what is the problem

OpenStudy (anonymous):

o.k, so first look at ur negative

OpenStudy (dumbcow):

well subtracting a negative is same as adding, 2 neg make a pos find LCM of 35 and 40 35 = 7*5 40=8*5 LCm = 5*8*7 = 280 35*8 = 280 40*7 = 280 -8/280 + 63/280 =55/280 =11/56

OpenStudy (anonymous):

yeah....what the cow said.....mooooooo

OpenStudy (anonymous):

11/56

OpenStudy (anonymous):

alright thank you

OpenStudy (anonymous):

do you understand it?

OpenStudy (anonymous):

I kind of do

OpenStudy (anonymous):

thats good, now practice, practice, practice

OpenStudy (anonymous):

yeah I will

OpenStudy (anonymous):

good, one medal for you!!!!!

OpenStudy (anonymous):

monkey out......

OpenStudy (anonymous):

oh thanks for the metal

OpenStudy (anonymous):

I just dont understand where the -8 came from in the above equation

OpenStudy (anonymous):

ok, so the only way you will have to form the same LCD is by mult. (8/8), its just one, but it allows you to share a common denominator so that you can add/subtract the two fractions.

OpenStudy (anonymous):

oh okay I did that with the -1/35 and got 8 over 280

OpenStudy (anonymous):

yeah, so u will get -8/280

OpenStudy (anonymous):

oh okay I understand it now thanks a lot man